# FairMT-Bench: Benchmarking Fairness for Multi-turn Dialogue in Conversational LLMs

### v1

@@ 4 Related Work
Recent research has revealed that LLMs tend to inherit biases from their pre-training data (Navigli et al., 2023). To explore and address the fairness issues in LLMs, numerous efforts have been made to evaluate these models.  Parrish et al. (2021) and  Li et al. (2020) created datasets and frameworks to evaluate social biases in question-answering models, focusing on contexts with varying amounts of information and bias-laden prompts. Similarly,  Wan et al. (2023),  Sun et al. (2024a), and  Wang et al. (2024) developed automated and template-based approaches for identifying and measuring social biases in conversational AI systems, covering a range of bias types and demographics across different tasks. While these works provide robust benchmarks for fairness evaluation, they largely concentrate on single-turn dialogues, neglecting the complexities of multi-turn interactions such as bias accumulation and contextual interference. Although many models demonstrate strong performance on these fairness benchmarks, current evaluations primarily focus on single-turn dialogues, overlooking the complexities introduced in multi-turn dialogues.
@@
Multi-turn dialogues are closer to real-world scenarios and play a crucial role in enhancing user experience. Some research force on multi-turn dialogue capability have shown that, in multi-turn dialogues, LLMs exhibit significant challenges in maintaining logical consistency, topic adaptability, memory retention, and contextual understanding (Duan et al., 2023). For example, as the conversation history grows, LLMs experience noticeable declines in memory retention, coherence, and adaptability (Bai et al., 2024). Additionally, when dealing with multi-turn instructions involving pronouns and ellipses, LLMs demonstrate considerable difficulties in understanding (Sun et al., 2024b). The increased complexity in multi-turn dialogues can also expose vulnerabilities in LLMs’ safety mechanisms that remain undetected in single-turn settings (Li et al., 2024; Yang et al., 2024). For example, toxicity can accumulate in seemingly harmless contexts (Zhou et al., 2024; Chen et al., 2023).
@@ 5 Conclusion
In this paper, we introduce the first benchmark, FairMT-Bench, for evaluating model fairness in multi-turn dialogues. We developed a comprehensive taxonomy to guide task design and generated the FairMT-10K dataset, covering nearly all bias types and attributes commonly addressed in fairness evaluation, representing comprehensiveness. Testing six LLMs on this dataset showed models are more prone to biased responses in multi-turn dialogues, with performance varying across tasks due to differences in multi-turn capabilities. We also created the more challenging FairMT-1K dataset, testing 15 state-of-the-art LLMs, all of which exhibited high bias. Our benchmark highlights the need for improved fairness alignment in complex multi-turn dialogue scenarios.
@@
This research on LLM fairness employs publicly available datasets, ensuring compliance with privacy regulations and anonymization where required. Our objective is to advocate for the responsible and fair utilization of LLMs, thereby enhancing their trustworthiness and reliability, while promoting ethical AI development. This study does not involve human subjects or breach any legal compliance standards.
@@ A.1 Task Design
In this section, we present the specific design of each task in the Fair-MT Bench and provide templates for generating multi-turn prompts.
@@ A.4.1 Instructions to Crowdsourcers
A.5 Results of Human Evaluation
Three annotators labeled the sampled data based on our Annotation Guideline. When there was disagreement among the three annotators, we used a voting mechanism to select the result agreed upon by two annotators. In the end, the consistency between the human annotations and GPT-4 annotations was 87.5%. In the sampled data, the bias rate in human annotations was 30.27%, while the bias rate in GPT-4 annotations was 36.61%. We conducted a detailed analysis of the samples with inconsistent annotations and found that when the model’s response presented a bias by explaining it through certain objective phenomena or the model adds a seemingly fair explanation to a biased answer., human annotators tended to label such responses as non-biased, while GPT-4 considered any attempt to associate a group with a particular stereotype, regardless of the reason, as biased. An example is shown in Figure 18.
Figure 18: An example of a discrepancy between human annotations and GPT annotations.
@@ B.1 EVALUATION Result Evaluated by GPT-4
We presented a heatmap to visualize the evaluation results of GPT-4 as shown in Figure 19. Compared to the previous Llama-Guard results, GPT-4’s evaluation results are more dispersed, but certain trends can still be observed. For example, in the Stereotype and Toxicity results, a concentrated dark area appears in the upper-right corner of the heatmap. This indicates that models with stronger language understanding capabilities, such as ChatGPT and Llama-3.2-8B-Instruct, are more sensitive to interactions and interference from user instructions.
Figure 19: Distribution of bias rates of models evaluated by GPT-4 on different tasks.
Table 4: The bias ratio of the model on different tasks evaluated by Llama-Guard-3. The bias rate indicates the proportion of biased answers in the answers generated by the model in the last turn to all answers.
Stereotype
ChatGPT 0.15% 0.55% 0.00% 28.05% 1.15% 0.00% 4.98%
Llama-3.1-8B-Instruct 0.29% 1.34% 0.36% 39.96% 20.96% 0.96% 10.64%
Gemma-7b-it 3.44% 0.28% 1.20% 41.15% 5.14% 0.14% 8.56%
Llama-2-7b-chat 2.15% 4.39% 59.57% 8.32% 2.23% 2.28% 13.16%
Llama-2-13b-chat 0.29% 6.31% 42.96% 5.51% 0.49% 2.07% 9.60%
Toxicity
ChatGPT 0.00% 1.34% 0.00% 46.57% 12.27% 0.00% 10.03%
Llama-3.1-8B-Instruct 0.00% 1.09% 0.36% 25.99% 36.17% 0.11% 10.62%
Gemma-7b-it 2.98% 2.91% 1.20% 23.08% 2.29% 0.00% 5.41%
Llama-2-7b-chat 2.21% 33.71% 26.57% 2.91% 0.00% 46.23% 18.60%
Llama-2-13b-chat 2.63% 28.16% 37.96% 7.48% 0.00% 49.00% 20.87%
@@ B.2 Case Study
Figure 20: A case study of Anaphora Ellipsis.
Figure 21: A case study of Scattered Questions.
@@ B.4 EVALUATION RESULTS ON DIFFERENT GROUPS
Here we show a color bar zoomed-in version of differences in model performance on different bias attributions in Figure 22.
@@ Appendix C Related Works: LLM Fairness evaluation
More recent research has focused on evaluating fairness in open-text conversations. Parrish et al. (2021) and Li et al. (2020) created datasets and frameworks to evaluate social biases in question-answering models, focusing on contexts with varying amounts of information and bias-laden prompts. Similarly, Wan et al. (2023), Sun et al. (2024a), and Wang et al. (2024) developed automated and template-based approaches for identifying and measuring social biases in conversational AI systems, covering a range of bias types and demographics across different tasks. While these works provide robust benchmarks for fairness evaluation, they largely concentrate on single-turn dialogues, neglecting the complexities of multi-turn interactions such as bias accumulation and contextual interference. Although many models demonstrate strong performance on these fairness benchmarks, current evaluations primarily focus on single-turn dialogues, overlooking the complexities introduced in multi-turn dialogues.

### v2

@@ 4 Related Work
Recent research has revealed that LLMs tend to inherit biases from their pre-training data (Navigli et al., 2023). To explore and address the fairness issues in LLMs, numerous efforts have been made to evaluate these models.  Parrish et al. (2021) and  Li et al. (2020) created datasets and frameworks to evaluate social biases in question-answering models. Similarly,  Wan et al. (2023),  Sun et al. (2024a), and  Wang et al. (2024) developed automated and template-based approaches for identifying and measuring social biases in conversational AI systems. While these works provide robust benchmarks for fairness evaluation, they largely concentrate on single-turn dialogues, neglecting the complexities of multi-turn interactions such as bias accumulation and contextual interference. A detailed comparison with exiting works is provided in Appendix C.
@@
Multi-turn dialogues are closer to real-world scenarios and play a crucial role in enhancing user experience. Some research have shown that, LLMs exhibit significant challenges in in multi-turn dialogues (Bai et al., 2024; Duan et al., 2023). Additionally, when dealing with multi-turn instructions involving pronouns and ellipses, LLMs demonstrate considerable difficulties in understanding (Sun et al., 2024b). The increased complexity in multi-turn dialogues can also expose vulnerabilities in LLMs’ safety mechanisms that remain undetected in single-turn settings (Li et al., 2024; Yang et al., 2024; Zhou et al., 2024; Chen et al., 2023).
@@ 5 Conclusion
In this paper, we introduce the benchmark, FairMT-Bench, for evaluating fairness of LLMs in multi-turn dialogues. We develop a comprehensive taxonomy to guide task design and generate the FairMT-10K dataset, covering nearly all bias types and attributes commonly address in fairness evaluation. Testing 6 LLMs on this dataset shows that LLMs are more prone to biased responses in multi-turn dialogues, with performance varying across tasks due to differences in multi-turn capabilities. We also distill a more challenging FairMT-1K dataset, testing 15 state-of-the-art LLMs, all of which exhibited high bias. Our benchmark underscores the need for improved fairness alignment in multi-turn scenarios, based on the LLMs multi-turn dialogue capabilities and fairness performance.
@@
This study on the fairness of large language models (LLMs) uses publicly available datasets as data sources, ensuring compliance with privacy regulations and anonymizing data when necessary. Although the data may contain some toxic or stereotypes, the purpose of using this data is to test whether the model can effectively identify these issues, not to propagate toxicity or stereotypes. Our goal is to advocate for the responsible and fair use of LLMs to enhance their credibility and reliability, and to promote the development of ethical artificial intelligence. We have made the dataset and code implementation public to ensure transparency throughout the research process.
In terms of the methods and basis for constructing the dataset, this study draws on existing research about the capabilities of LLMs in multi-turn dialogues and techniques for bypassing restrictions. During manual verification, we provided professional training and detailed guidelines to annotators. Each piece of data was independently reviewed by three annotators, with final results aggregated through a voting process to minimize individual biases impacting the annotation outcomes. The research process adheres strictly to all legal and regulatory standards. Specific details of the annotation process can be found in the appendix. We ensure that the methods used to construct the dataset and the content of the dataset comply with ethical standards.
@@
Acknowledgement
This work is supported by the National Natural Science Foundation of China (Grant No. 62476241), the Natural Science Foundation of Zhejiang Province, China (Grant No. LZ23F020008), and the Zhejiang University-Angelalign Inc. R&D Center for Intelligent Healthcare.
@@ A.1 Task Design
In this section, we present the specific design of each task in the Fair-MT Bench and provide templates for generating multi-turn prompts. The distribution of tasks in the final FairMT-10K dataset is shown in Table 4.
@@
Table 4: The number distribution of each task on the FairMT-10K dataset
Bias Type Scattered Questions Anaphora Ellipsis Jailbreak Tips Interference Misinformation Negative Feedback Theme Variations
Stereotype 1356 1211 841 1356 1356 1356
Toxicity 481 459 298 481 481 481
Total 1837 1670 1139 1837 1837 1837
@@ A.4.1 Instructions to Crowdsourcers
A.4.2 Results of Human Evaluation
Three annotators labeled the sampled data based on our Annotation Guideline. We calculated the consistency between the human assessment results and those of GPT-4. The results of this consistency assessment are presented in Table 5. Overall, human validation and GPT-4 evaluation results show high accuracy and recall values, indicating strong consistency between the two. GPT-4 rarely exhibits cases of missed judgments. However, in certain tasks, GPT-4’s evaluations demonstrate a tendency toward over-protection. We conducted a detailed analysis of the samples with inconsistent annotations and found that when the model’s response presented a bias by explaining it through certain objective phenomena or the model adds a seemingly fair explanation to a biased answer., human annotators tended to label such responses as non-biased, while GPT-4 considered any attempt to associate a group with a particular stereotype, regardless of the reason, as biased. An example is shown in Figure 18.
A.5 The cost of using the FairMT Bench.
Time cost
The generation speed varies among different models. In this analysis, we take Llama-2-7b-chat model as an example. Some API-called models, such as ChatGPT, generally produce faster results. We configure the model to generate text with “max_new_tokens” set to 150 and maintain the batch size and precision as 1 and FP32, respectively. The generation is performed on a single NVIDIA H100 GPU. Under this setup, generating 10K samples of 5-turn multi-turn dialogue responses takes about 72.5 H100 GPU hours.
Economic cost
First, we discuss the costs associated with using GPT-4 as an evaluator. Specifically, we use GPT-4 Turbo to evaluate the generated content of the test models. The current price of GPT-4 Turbo is \$0.01 per 1,000 input tokens and \$0.03 per 1000 output tokens. In our setup, GPT-4 evaluates bias in the final turn of the dialogue by processing the historical dialogue as context and providing simple outputs: “Yes, it is biased” or “No, it is not biased”. We have estimated the cost of evaluating each dialogue set to be approximately 0.017 USD. In addition, we also use the open-source model LlamaGuard-3 as another evaluation tool. While LlamaGuard-3 is less sensitive in detecting implicit biases compared to GPT-4, its assessment results and trends are generally consistent with those of GPT-4. LlamaGuard-3 serves as a cost-effective alternative when GPT-4 API is unavailable or to minimize expenses. Furthermore, to enhance evaluation efficiency and reduce costs, we have curated 1,000 particularly challenging samples from the FairMT-10K dataset into a new subset, FairMT-1K.
Table 5: Consistency between new Human Validation and GPT-4 Results.
Task Acc Recall Precision F1
Scattered Questions 0.9569 0.9394 0.9254 0.9323
Anaphora Ellipsis 0.9646 0.9490 0.9789 0.9637
Jailbreak Tips 0.9020 1.0000 0.4000 0.5714
Interference Misinformation 0.9236 0.9655 0.8615 0.9106
Negative Feedback 1.0000 1.0000 1.0000 1.0000
Theme Variations 0.9268 0.9841 0.8493 0.9118
Total 0.9569 0.9646 0.9160 0.9397
@@ B.1 EVALUATION Result Evaluated by GPT-4
We presented a heatmap to visualize the evaluation results of GPT-4 as shown in Figure 19. Compared to the previous Llama-Guard results, GPT-4’s evaluation results are more dispersed, but certain trends can still be observed. For example, in the Stereotype and Toxicity results, a concentrated dark area appears in the upper-right corner of the heatmap. This indicates that models with stronger language understanding capabilities, such as ChatGPT and Llama-3.1-8B-Instruct, are more sensitive to interactions and interference from user instructions.
Additionally, We tested the multi-turn dialogue capabilities of the six models on MT-Bench 101  (Bai et al., 2024), a benchmark for fine-grained evaluation of model multi-turn dialogue capabilities.
In our FairMT-Bench, we designed tasks to assess the trade-off between fairness and model utility. In conjunction with our tasks for assessing model fairness, we focused on the trade-off between fairness and model utility, observing whether the model begins to refuse responses due to keywords associated with certain social groups or attributes without any biased intent from the user, sacrificing model utility to ensure fairness. In our task design, the Fixed Format task involves asking the model to answer unbiased questions related to a social group in a fixed format for the first four turns, and then posing a biased question related to the same social group in the same format in the fifth turn. This task design considers the model’s ability to balance performance and fairness under clear instructions and requirements, and also examines the model’s flexibility in changing its response strategy when the dialogue content undergoes subtle changes across multiple turns. Ideally, the model should normally answer according to the format in the first four turns and refuse to answer while pointing out the bias in the prompt in the last turn of dialogue. Here, we examine whether the model exhibits over-protection by generally refusing to respond upon detecting sensitive words related to a particular group, specifically if the model displays refusal behavior in the first four turns.
The test results from MT-Bench 101 (Bai et al., 2024) and the model’s over-refusal are shown in Table 6. Overall, models with stronger multi-turn dialogue capabilities tend to exhibit better fairness performance in multi-turn dialogues (as analyzed from the Avg. results). Specifically, we found that models with stronger Adaptability and Interactivity are more susceptible to user guidance and interference, which can alter their fairness strategies during multi-turn dialogues. We think this may be because the fairness alignment of these models has not adequately accounted for multi-turn adversarial prompts.
Table 6: The evaluation results of the models’ multi-turn dialogue capabilities based on MT-Bench 101, as well as the over-refusal ratio of the models on the Fixed Format tasks, are presented.
MT-Capability
Model Perceptivity Adaptability Interactivity Avg. Over Refusual
ChatGPT 8.73 7.55 7.40 7.99 0.00%
Llama-2-7b-chat 7.70 6.00 5.17 6.29 29.89%
Llama-2-13b-chat 8.47 6.39 6.15 7.15 23.10%
Llama-3.1-8B-Instruct 4.26 2.37 3.40 3.34 12.78%
Mistral-7B-Instruct-v0.3 7.85 6.82 6.00 6.89 21.40%
Gemma-7B-it 8.93 7.03 5.26 7.07 26.03%
Figure 19: Distribution of bias ratio of models evaluated by GPT-4 on different tasks.
@@ B.3 Comparison of Performance between Single and Multi-Turn
Table 8: The bias ratio of the model on different tasks evaluated by Llama-Guard-3. The bias ratio indicates the proportion of biased answers in the answers generated by the model in the last turn to all answers.
Stereotype
ChatGPT 0.15% 0.55% 0.00% 28.05% 1.15% 0.00% 4.98%
Llama-3.1-8B-Instruct 0.29% 1.34% 0.36% 39.96% 20.96% 0.96% 10.64%
Gemma-7b-it 3.44% 0.28% 1.20% 41.15% 5.14% 0.14% 8.56%
Llama-2-7b-chat 2.15% 4.39% 59.57% 8.32% 2.23% 2.28% 13.16%
Llama-2-13b-chat 0.29% 6.31% 42.96% 5.51% 0.49% 2.07% 9.60%
Toxicity
ChatGPT 0.00% 1.34% 0.00% 46.57% 12.27% 0.00% 10.03%
Llama-3.1-8B-Instruct 0.00% 1.09% 0.36% 25.99% 36.17% 0.11% 10.62%
Gemma-7b-it 2.98% 2.91% 1.20% 23.08% 2.29% 0.00% 5.41%
Llama-2-7b-chat 2.21% 33.71% 26.57% 2.91% 0.00% 46.23% 18.60%
Llama-2-13b-chat 2.63% 28.16% 37.96% 7.48% 0.00% 49.00% 20.87%
@@ B.4 EVALUATION RESULTS ON DIFFERENT GROUPS
From the model’s performance across different social attributes, certain attributes pose significant challenges for all models. Here, we provide a brief discussion of this phenomenon. As  Gallegos et al. (2024) mentioned that, the primary source of bias in large language models (LLMs) is the training data. Similarly, we believe that the effectiveness of fairness alignment in models largely depends on the data used for alignment. Specifically, the most notable instance in the FairMT-Bench test results appears in the category of gender stereotypes. We selected instances where the model’s responses exhibited bias for detailed analysis and discussion of common issues.
In the case of gender stereotypes, biased responses primarily manifest as very subtle biases, such as seemingly conventional and normative implicit biases against women. Examples include associations with domesticity and compliance, which are generally not recognized by models as biases that would trigger protective mechanisms. A case study is shown in Figure 20. Here we show a color bar zoomed-in version of differences in model performance on different bias attributes in Figure 21.
Figure 20: A case study on a model’s difficulty in avoiding overly implicit biases.
B.5 FairMT-1K construction process
To reduce the cost of evaluating fairness in multi-turn dialogues and enhance the efficiency of these evaluations, we selected a more challenging dataset based on the results of six models tested on the FairMT-10K dataset. We assume that the more models that exhibit biased responses to a data point, the more challenging that data point is for the models. Therefore, we selected the composition of the FairMT-1K dataset based on the number of times a model exhibited bias on a particular data point, independent of the model’s fairness performance. For each task, we integrated the evaluation results of six models across two types of biases and counted the number of models that exhibited bias in the final turn of each dialogue group. We considered a dialogue group as challenging if more models produced biased responses in that group. We ranked all samples based on the number of models that provided biased responses and selected the top 170 groups with the most biased responses to be included in the FairMT-1K dataset.
B.6 Bias Mitigation by Fairness Alignment
We employed commonly used alignment paradigms DPO (Rafailov et al., 2024) for debiasing. We constructed positive and negative data pairs for fairness multi-turn alignment based on the evaluation results on the FairMT-10K dataset. We use the first four turns of historical dialogue as the “conversation”, with the unbiased response from the model serving as the “chosen” answer and the biased response as the “rejected” answer. The base model is Llama-2-7b-chat. After debiasing, we assess the model’s fairness, multi-turn dialogue capabilities, and the proportion of over-protection after alignment. The details of capabilities and over-protection evaluation. Specifically, we evaluate the model’s multi-turn dialogue capabilities based on MT-Bench 101 (Bai et al., 2024) and calculate the proportion of over-protection by determining the model’s refusal ratio in the Fixed Format tasks for the first four turns where the questions are entirely unbiased. To further investigate the performance in fairness alignment, we have also compared the results on a single-turn alignment dataset using only the last turn conversation of FairMT-10K dataset. The results are shown in Table 9.
From the experimental results, the model’s bias ratio significantly decreased after alignment, and the performance of fairness under both single and multi-turn alignment paradigms was essentially consistent, suggesting that improvements in fairness primarily arose from the model fitting to the dataset used in this work. However, both single and multi-turn aligned models showed a decline in multi-turn dialogue language modeling capabilities, and an increase in the proportion of over-refusal, indicating that the model sacrificed performance to a large extent to enhance fairness, without achieving a good trade-off. However, it can also be seen that multi-turn alignment had a smaller impact on the model’s language modeling capabilities compared to the currently prevalent single-turn safety alignment, showing certain advantages of multi-turn alignment. Therefore, designing more high-quality multi-turn fairness alignment datasets and techniques to achieve a trade-off between fairness and quality remains an unresolved issue.
Table 9: Fairness and quality performance after model alignment.
Metrics Base Model
ST-Align-
on-FairMT
MT-Align-
on-FairMT
Bias Ratios ↓ Scattered Questions 94.55% 0.00% 0.00%
Anaphora Ellipsis 43.03% 5.45% 4.85%
Jailbreak Tips 69.09% 0.00% 0.00%
Interference Misinformation 73.94% 11.52% 10.91%
Negative Feedback 88.48% 1.21% 1.21%
Theme Variations 89.09% 49.09% 55.76%
Avg. 76.36% 11.21% 12.12%
MT-Capability ↑ Perceptivity 7.70 6.35 6.77
Adaptability 6.00 6.78 6.01
Interactivity 5.17 2.60 4.43
Avg. 6.29 5.24 5.73
Over Refusual ↓ 12.36% 23.03% 21.82%
@@ Appendix C Related Works: LLM Fairness evaluation
More recent research has focused on evaluating fairness in open-text conversations. Parrish et al. (2021) and Li et al. (2020) created datasets and frameworks to evaluate social biases in question-answering models, focusing on contexts with varying amounts of information and bias-laden prompts. Similarly, Wan et al. (2023), Sun et al. (2024a), and Wang et al. (2024) developed automated and template-based approaches for identifying and measuring social biases in conversational AI systems, covering a range of bias types and demographics across different tasks. While these works provide robust benchmarks for fairness evaluation, they largely concentrate on single-turn dialogues, neglecting the complexities of multi-turn interactions such as bias accumulation and contextual interference. Although many models demonstrate strong performance on these fairness benchmarks, current evaluations primarily focus on single-turn dialogues, overlooking the complexities introduced in multi-turn dialogues. Furthermore, as shown in Table 10, we compare our work with previous related work in terms of tasks, social groups, and types of bias, demonstrating that FairMT-Bench is a novel and comprehensive benchmark.
Similar to fairness, the domain of safety has also seen the emergence of many comparable evaluation efforts like S-Eval (Yuan et al., 2024) and CoSafe (Yu et al., 2024), which are gradually building multi-round, comprehensive evaluation frameworks. Drawing on experience from the safety domain, we have expanded to a fine-grained, multi-turn-capability-based benchmark, which is also rare in the field of safety.
Figure 22: A case study of Anaphora Ellipsis.
Figure 23: A case study of Scattered Questions.
Table 10: Comparison between FairMT-Bench with other fairness evaluation works. $\checkmark$ indicates the inclusion of corresponding information, - indicates that the related information is not discussed in the paper, and * represents a multiplication sign.
Benchmark Task Social Group Bias type # Num
Gender Race Age Religion Disabled Appearance Stereotype Toxicity
StereoSet (Nadeem et al., 2020) Embedding $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 16,995
Crows-Pairs (Nangia et al., 2020) Embedding $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 1,508
Redditbias (Barikeri et al., 2021) Embedding $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 11,873
Jigsaw (cjadams, 2019) Embedding $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 50,000
BBQ (Parrish et al., 2021) Conversation $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 58,492
UnQover (Li et al., 2020) Conversation $\checkmark$ 30*
BiasAsker (Wan et al., 2023) Conversation $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 841 *8,110
BOLD (Dhamala et al., 2021) Text Completion $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 23,679
HONEST (Nozza et al., 2021) Text Completion $\checkmark$ $\checkmark$ 420
HolisticBias (Liang et al., 2022) Conversation $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 600*13*6
RealToxicityPrompt (Gehman et al., 2020) Conversation $\checkmark$ 100,000
TrustLLM (Sun et al., 2024a) Conversation $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ -
CEB (Wang et al., 2024) Five Tasks [^1] Specifically divided into Indirect Tasks and Direct Tasks. Indirect Tasks include Continuation, Conversation and Classification, while Direct Tasks include Recognition and Selection. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 11,004
Ours (FairMT) Multi-Turn Conv. $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ $\checkmark$ 10,157